### v1

# MFCPL: Multimodal Federated Cross Prototype Learning for Missing Modalities and Data Heterogeneity

Abstract
Multimodal federated learning (MFL) has emerged as a decentralized machine learning paradigm, allowing multiple clients with different modalities to collaborate on training a machine learning model across diverse data sources without sharing their private data. However, challenges, such as data heterogeneity and severely missing modalities, pose crucial hindrances to the robustness of MFL, significantly impacting the performance of global model. The absence of a modality introduces misalignment during the local training phase, stemming from zero-filling in the case of clients with missing modalities. Consequently, achieving robust generalization in global model becomes imperative, especially when dealing with clients that have incomplete data. In this paper, we propose Multimodal Federated Cross Prototype Learning (MFCPL), a novel approach for MFL under severely missing modalities by conducting the complete prototypes to provide diverse modality knowledge in modality-shared level with the cross-modal regularization and modality-specific level with cross-modal contrastive mechanism. Additionally, our approach introduces the cross-modal alignment to provide regularization for modality-specific features, thereby enhancing overall performance, particularly in scenarios involving severely missing modalities. Through extensive experiments on three multimodal datasets, we demonstrate the effectiveness of MFCPL in mitigating these challenges and improving the overall performance.
1 Introduction
Figure 1: Problem Illustration of Multimodal Federated Learning. We present two challenges of MFL which are a) Missing Modality and b) Data Heterogeneity.
Federated learning (FL) McMahan et al. (2017); Kairouz et al. (2021); Tan et al. (2022a) has emerged as a promising distributed machine learning paradigm, allowing multiple clients to collaboratively train models without leaking private data. Despite promising outcomes observed across various domains Rieke et al. (2020); Nguyen et al. (2021), most FL methods are tailored for scenarios with clients processing unimodal data, potentially constraining their applicability in real-world settings. However, in the era of data proliferation, users have the distinctive opportunity to harness many data types, unlocking new dimensions to explore the potential of multimodal data that aligns more closely with real-world applications Caesar et al. (2020); Fu et al. (2023). Therefore, with the necessity of handling multimodal data while preserving privacy, multimodal federated learning (MFL) has emerged as a promising paradigm Xiong et al. (2022); Zhao et al. (2022). In traditional FL scenarios, data heterogeneity Li et al. (2020); Deng et al. (2020), where distributed data is non-IID (identically and independently distributed), has become a crucial challenge, leading to performance degradation. Despite numerous studies in FL domain Luo et al. (2021); Mendieta et al. (2022) aimed at mitigating data heterogeneity, their predominant focus has been on addressing this challenge in applications involving unimodal data, rather than extending these solutions to the context of MFL. Additionally, in practical applications, it is essential to recognize that real-world scenarios may not always allow the utilization of all available modalities. Challenges such as sensor disconnections, different hardware platforms, or other operational issues can impede the efficient integration of diverse data streams, impacting the model’s robustness. While data heterogeneity presents its own set of challenges, the situation is compounded by the significant impact of severely missing modalities, leading to a further degradation in the model’s performance within the context of MFL Chen and Zhang (2022); Feng et al. (2023). Consequently, we pose a research question: How can we effectively address both data heterogeneity and missing modality challenges in MFL?. In Figure 1, we illustrate the concepts of data heterogeneity and missing modalities.
In the context of MFL, when confronted with significant missing modalities, a common practice is to impute the missing data with zeros. However, this zero-filling strategy introduces potential misalignments or undesired effects, adversely impacting the model performance in multimodal learning. To enhance representation learning in MFL under conditions of missing modalities, an effective guiding method during training is crucial. Recently, prototype learning has been introduced as an effective information carrier between the server and clients in FL Tan et al. (2022c); Qiao et al. (2023). A prototype is defined as the mean value of features from the same class and is represented as a vector. Several studies in FL have leveraged prototypes to tackle issues such as data heterogeneity Tan et al. (2022b) or domain shift Huang et al. (2023). Nonetheless, prior research on FL with prototypes has solely focused on scenarios involving unimodal data. Building upon the inspiration derived from prototype learning, we aim to extend its application to address the challenge of missing modalities in MFL. To tackle the challenges of both data heterogeneity and the existence of missing modalities, we derive the complete prototype by aggregating local prototypes from various clients, including those with both missing and full modalities, utilizing the shared projection head. This ensures that the features of each class encapsulate knowledge from diverse modalities. Additionally, the complete prototype serves as guidance for clients experiencing missing modalities.
In this paper, we propose Multimodal Federated Cross Prototype Learning (MFCPL), which consists of three main components: Cross-Modal Prototypes Regularization (CMPR), Cross-Modal Prototypes Contrastive (CMPC) and Cross-Modal Alignment (CMA). Drawing inspiration from the success of prototypes in addressing data heterogeneity in FL Tan et al. (2022b); Mu et al. (2023), we introduce a novel complete prototype to effectively handle missing modalities in MFL. First, we propose Cross-Modal Prototypes Regularization (CMPR) by shortening the distance between complete prototype and the local modality-shared representation, bridging the gap between local and global models. Through this process, diverse modality knowledge is effectively transferred from complete prototypes to those clients with missing modalities, thereby contributing to the robustness of the learning system under conditions of missing modalities. Second, we introduce Cross-Modal Prototypes Contrastive (CMPC), leveraging complete prototypes to facilitate cross-modal contrastive learning. CMPC encourages local modality-specific representations to be close to the complete prototypes within the same semantic class and pushes them away from complete prototypes of different classes. Cross-modal contrastive learning involves aligning modality-specific representations and complete prototypes into the same latent space through a projection layer. We can provide robust and rich semantic representations suitable for clients with missing modalities by aligning the complete prototype with modality-specific representations. Finally, we introduce Cross-Modal Alignment (CMA), which aligns the embedding features of the missing modality with those of the existing modality through a projection layer. By incorporating this, we mitigate the negative impact of missing modality and improve the generalization performance of the global model. These components collectively empower our proposed framework as a robust method to tackle severe missing modalities in heterogeneous MFL effectively. Our primary contributions are:
  • To the best of our knowledge, we are the first in heterogeneous MFL to integrate the concept of complete prototype for addressing the missing modality challenge. Leveraging the prototype perspective, we enhance performance by calibrating the local training process with diverse modality knowledge, particularly benefiting clients dealing with missing data.
  • We propose a novel multimodal federated learning framework, MFCPL, to learn the global model with severe missing modality. Building on the success of prototypes in FL, we introduce complete prototypes aligned at two levels, offering rich semantic information from multiple modalities: modality-specific representation with CMPC and modality-shared representation with CMPR through a projection layer. Additionally, we address the negative impact associated with missing modality by introducing the loss function for CMA.
  • We comprehensively evaluate the performance of MFCPL through extensive experiments on three multimodal federated datasets, demonstrating its superiority over other baseline methods. Complemented by ablation studies, the promising results validate the effectiveness of each component.
2 Related Work
2.1 Multimodal Federated Learning
Numerous research works in multimodal federated learning (MFL) Xiong et al. (2022); Zhao et al. (2022); Yu et al. (2023), have explored diverse techniques to facilitate the decentralized training of machine learning models across multiple data sources. For instance, Xiong et al. Xiong et al. (2022) introduced a co-attention layer to merge representations from various modalities, obtaining global features to train personalized models for clients. Addressing labeled data constraints in MFL, Zhao et al. Zhao et al. (2022) proposed a semi-supervised framework within the FedAvg mechanism McMahan et al. (2017). CreamFL Yu et al. (2023) introduced a knowledge distillation-based MFL framework using a contrastive method for representation aggregation leveraging the public dataset. However, these methods do not consider the missing modality scenario in the system, limiting their practical applicability in real-world settings. To address the issue of missing modalities in MFL, Chen and Zhang Chen and Zhang (2022) employed a dynamic graph structure to capture correlations among different types of clients. Recently, FedMultimodal Feng et al. (2023) introduced the first benchmark in MFL that evaluates MFL under different types of data corruption: missing modalities, missing labels, and erroneous labels. In this paper, to address the missing modality problem in heterogeneous MFL, we introduce the novel concept of leveraging a complete prototype to operate with local representations at two levels: modality-specific representations and modality-shared representation through a projection layer.
2.2 Incomplete Multimodal Learning
Exploring incomplete data, such as missing modalities in multimodal learning, is crucial while ensuring robustness in real-world scenarios with missing modalities. Instances of missing modalities are prevalent in practical situations, often stemming from hardware constraints or storage limitations. Various approaches, such as data imputation Tran et al. (2017) or data reconstruction Pham et al. (2019); Zhao et al. (2021), have been proposed to address these challenges by filling in missing modalities based on existing ones. For instance,  Tran et al. Tran et al. (2017) proposed a cascaded residual autoencoder (CRA) to impute missing modalities. Additionally, Dong et al. Dong et al. (2023) introduced a cross-modal translation method to replace null entries in the embedding for missing modality generalization during the inference stage. Another approach involves utilizing a knowledge transfer mechanism to convey information from modalities with available data to those with missing data Garcia et al. (2018); Luo et al. (2018)Poklukar et al. Poklukar et al. (2022) introduced the Geometric Multimodal Contrastive (GMC) framework, which geometrically aligns complete and modality-specific representations to address missing modality information across various tasks. In this study, we employ the definition of a complete prototype to provide diverse modality knowledge for clients with missing modalities, thereby enhancing the local training process for these clients.
2.3 Prototype Learning
The prototype is computed as the mean feature vectors within the same class Snell et al. (2017) and has found applications in various research domains, including few-shot learning Xu et al. (2020); Zhang et al. (2021), and unsupervised learning Ye et al. (2019); Guo et al. (2022). In FL, prototypes have been employed in multiple studies to address data heterogeneity Tan et al. (2022b, c); Huang et al. (2023); Qiao et al. (2023). FedProto Tan et al. (2022b) was among the pioneering works to apply prototypes in FL, introducing a communication approach between clients and the server using prototypes instead of parameter exchange to learn personalized models. Additionally,  Qiao et al. Qiao et al. (2023) proposed the multi-prototype approach with the contrastive learning strategy to address the feature heterogeneity. Recently, FPL Huang et al. (2023) proposed the cluster prototypes to tackle the domain shift in FL. However, the majority of prior works concentrate on scenarios where clients exclusively possess unimodal data. In the context of MFL under missing modalities, it is crucial to address the generalization across various types of clients with diverse modality absences. Our work illuminates the use of the complete prototype to accomplish this objective at two levels: modality-specific and modality-shared representations.
3 Methodology
3.1 Preliminaries
Multimodal federated learning (MFL) aims to facilitate collaborative training among multiple clients on a shared global model, leveraging multimodal data while ensuring the data privacy. In this paper, we consider that there are $N$ clients (indexed by $i$) with private multimodal data $D_{i}=\{{x}^{A}_{i},{x}^{B}_{i},{y}_{i}\}$, where $x^{A}_{i}$ and $x^{B}_{i}$ represent samples from modalities A and B, respectively, and $y_{i}$ denotes the label associated with each sample.
Figure 2: Illustration of MFCPL. Different client types upload their prototypes to server using projection head $g_{1}$. We provide diverse modality knowledge in modality-shared level with $\mathcal{L}_{CMPR}$ and modality-specific level with $\mathcal{L}_{CMPC}$ using projection head $g_{2}$. The alignment loss $\mathcal{L}_{CMA}$ is introduced to reduce the noise in with between projected modality-specific representations of two modalities using projection head $g_{2}$.
Client models share the identical architecture, comprising three modules: encoders $f_{m}$, self-attention fusion module $h_{\text{self-attention}}$, and classifier $g$. The encoders take input samples $x^{A}$ and $x^{B}$ and encode them into $n$-dimensional feature vectors $z_{A}=f_{A}(x^{A}),z_{B}=f_{B}(x^{B})\in\mathbb{R}^{n}$. Subsequently, a self-attention fusion module $h_{\text{self-attention}}$ integrates the representations from two modalities into the same feature space, denoted as $e=h_{\text{self-attention}}(z_{A},z_{B})\in\mathbb{R}^{u}$. Finally, the classifier $g$ maps the multimodal feature $e$ to the logits output $z_{cls}=g(e)\in\mathbb{R}^{I}$. Specifically, given the model parameters for the whole network as $\theta$, and $D$ is the sum of samples over all clients, we formulate the global objective:
$\operatorname*{argmin}_{\theta}{L}(\theta)$ $=\sum^{N}_{i=1}\frac{|D_{i}|}{|D|}{\mathcal{L}}_{i}(\theta_{i},D_{i}).$ (1)
where, the loss function $\mathcal{L}_{i}$ is the task training loss $\mathcal{L_{S}}=L(z_{cls},y)$ for $i^{th}$ client.
3.2 Multimodal Prototype Federated Learning
Prior research works in federated prototype learning Tan et al. (2022b); Mu et al. (2023) have focused on scenarios where clients possess unimodal data. In the context of MFL under missing modalities, multiple types of clients exist: those missing modality $A$, those missing modality $B$, and those with complete modalities $A$ and $B$. Simply aggregating unimodal prototypes based on modality may introduce a bias toward the dominant modality and cause the performance degradation in FL under severely missing modalities. This prompts us to leverage prototypes from all modalities to develop a generalized model for different types of clients.
Complete Prototypes.
To facilitate alignment in the local training process, we utilize the projection head $g_{1}$ to map the multimodal representations $e$ from different client types into the shared representation space $\mathbb{R}^{d}$:
$\{r_{A}=g_{1}(e_{A}),r_{B}=g_{1}(e_{B}),r_{AB}=g_{1}(e_{AB})\}\in\mathbb{R}^{d}.$ (2)
We define the $k^{th}$ class prototype from different types of clients as:
$p^{k}_{Ai}=\frac{1}{|S^{k}_{Ai}|}\sum_{i\in S^{k}_{Ai}}r_{Ai},$ (3)
$p^{k}_{Bi}=\frac{1}{|S^{k}_{Bi}|}\sum_{i\in S^{k}_{Bi}}r_{Bi},$ (4)
$p^{k}_{ABi}=\frac{1}{|S^{k}_{ABi}|}\sum_{i\in S^{k}_{ABi}}r_{ABi},$ (5)
where $S^{k}_{Ai}$, $S^{k}_{Bi}$ and $S^{k}_{ABi}$ are subsets of $D_{i}$ belonging to $k^{th}$ class of different client types, respectively. As complementary to unimodal prototypes, we initially introduce complete prototypes by averaging local prototypes from different types of clients in the shared representation space $\mathbb{R}^{d}$. Specifically, we formulate the complete prototypes on the server as follows:
$P^{k}=\frac{1}{N}\sum_{i\in N}\{p^{k}_{Ai},p^{k}_{Bi},p^{k}_{ABi}\}\in\mathbb{R}^{d},$ (6)
where $P^{k}$ denotes the complete prototypes belonging to class $k$. Specifically, our complete prototypes guide the local training process at two levels: modality-specific and modality-shared representations.
3.3 Multimodal Federated Cross Prototype Learning
To achieve a generalized model for different types of clients under the condition of missing modalities, we utilize complete prototypes to provide diverse modality knowledge, supporting local clients at two levels: modality-specific and modality-shared representations. As shown in Figure 2, our proposed method comprises three main components: Cross-Modal Prototypes Regularization (CMPR), Cross-Modal Prototypes Contrastive (CMPC), and Cross-Modal Alignment (CMA).
Cross-Modal Prototypes Regularization.
Similar to other traditional FL methods, MFL encounters the challenge of data heterogeneity, and further complicated by the issue of missing modalities. In response to these challenges, we introduce a regularization loss that narrows the gap between local and global models, simultaneously providing diverse modality knowledge for clients with missing modalities at the modality-shared representation level. Drawing inspiration from FedProto Tan et al. (2022b), we incorporate a regularization loss between local projected multimodal representations $r_{i}$ and complete prototypes $P^{k}$ within the same semantic class $k$. However, our approach differs from FedProto, as our regularization loss serves as an extension to address the challenges posed by missing modalities in the context of multimodal federated learning. To implement this regularization, we utilize $\ell_{2}$ distance and introduce the Cross-Modal Prototypes Regularization (CMPR) as follows:
$\mathcal{L}_{CMPR}=\sum_{k}\|r^{k}_{i}-P^{k}\|_{2}^{2}$ (7)
where $r^{k}_{i}$ is the local projected multimodal representations of $k$ semantic class.
Algorithm 1 MFCPL
Input: communication rounds T, local epochs R, number of local clients N, local dataset $D_{i}$.
Output: Final global model $\theta_{t}$
1:  Server Execution:
2:  for $t=0,\dots,T$ do
3:     for $i=0,\dots,N$ do
3:        $\theta^{i}_{t},p_{Ai},p_{Bi},p_{ABi}\leftarrow\textbf{LocalUpdate}(\theta_{t},P)$
[…]
5:  end for
6:  Client Execution:
7:  LocalUpdate($\theta_{t},P$):
8:  for $r=0,\dots,R$ do
9:     for each batch $\in D_{i}=\{{x}^{A}_{i},{x}^{B}_{i},{y}_{i}\}$ do
10:        $z_{Ai}=f_{A}(x^{A}_{i}),z_{Bi}=f_{B}(x^{B}_{i})$
11:        $e_{i}=h_{\text{self-attention}}(z_{Ai},z_{Bi}),r_{i}=g_{1}(e_{i})$
12:        $z^{\prime}_{Ai}=g_{2}(z_{Ai}),z^{\prime}_{Bi}=g_{2}(z_{Bi})$
13:        $\mathcal{L}_{CMPR}\leftarrow(r_{i},P)$ in Eq. 7
14:        $\mathcal{L}_{CMPC}\leftarrow(z^{\prime}_{Ai},z^{\prime}_{Bi},P)$ in Eq. 11
15:        $\mathcal{L}_{CMA}\leftarrow(z^{\prime}_{Ai},z^{\prime}_{Bi})$ in Eq. 12
16:        $\mathcal{L}_{S}\leftarrow(z_{cls}^{i},y_{i})$
17:        $\theta^{i}_{t}\leftarrow\theta^{i}_{t}-\eta\nabla\mathcal{L}$
18:     end for
19:  end for
20:  $p_{Ai}=[p^{1}_{Ai},p^{2}_{Ai},\dots,p^{k}_{Ai}]$
21:  $p_{Bi}=[p^{1}_{Bi},p^{2}_{Bi},\dots,p^{k}_{Bi}]$
22:  $p_{ABi}=[p^{1}_{ABi},p^{2}_{ABi},\dots,p^{k}_{ABi}]$
23:  return $\theta^{i}_{t},p_{Ai},p_{Bi},p_{ABi}$
Cross-Modal Prototypes Contrastive.
We envision that modality-specific representations in clients with missing modalities can capture rich and diverse modality knowledge from the complete prototype. The modality-specific features of local data instances should align closely with their respective complete prototypes within the same semantic class while exhibiting dissimilarity to complete prototypes associated with different semantic classes. Regarding the data instances $\{{x}^{A}_{i},{x}^{B}_{i},{y}_{i}\}$, we initially encode them using encoder $f_{m}$ to acquire the modality-specific feature vectors $z_{Ai}=f_{A}({x}^{A}_{i}),z_{Bi}=f_{B}({x}^{B}_{i})\in\mathbb{R}^{n}$. To facilitate alignment between the modality-specific features and complete prototypes, we employ the projection head $g_{2}$ to project the modality-specific feature vectors of each modality into the shared representation space:
$z^{\prime}_{Ai}=g_{2}(z_{Ai}),z^{\prime}_{Bi}=g_{2}(z_{Bi})\in\mathbb{R}^{d},$ (8)
Let $p$ be the corresponding complete prototype $p\in P$, $C^{k}=P-P^{k}$ represents the prototypes with different semantic class from the local instances. Subsequently, the Cross-Modal Prototypes Contrastive (CMPR) loss where all modalities are available is expressed as follows:
$\mathcal{L}_{ConA}=-\log\frac{\sum_{p\in P^{k}}\exp(s(z^{\prime}_{Ai},p)/\tau)}{\sum_{p\in P}\exp(s(z^{\prime}_{Ai},p)/\tau)}$ (9)
$\mathcal{L}_{ConB}=-\log\frac{\sum_{p\in P^{k}}\exp(s(z^{\prime}_{Bi},p)/\tau)}{\sum_{p\in P}\exp(s(z^{\prime}_{Bi},p)/\tau)}$ (10)
$\mathcal{L}_{CMPC}=\mathcal{L}_{ConA}+\mathcal{L}_{ConB}$ (11)
where $\mathrm{s}(u,v)=u^{\top}v/\lVert u\rVert\lVert v\rVert$ denotes the cosine similarity between the projected modality-specific feature and complete prototypes, and $\tau$ is the contrastive hyper-parameter. In instances where clients lack modality A or B, the contrastive loss with $z^{\prime}_{Ai}$ or $z^{\prime}_{Bi}$ will be excluded. The goal of Eq. (11) is to encourage the modality-specific representations from clients with missing modalities to closely align with the corresponding complete prototypes in the shared embedding space, thereby enhancing the generalization of these clients. In this component, we utilize the projection head to facilitate alignment between feature vectors at different levels: the complete prototype and modality-specific representations. The effectiveness of the projection head is further discussed in the Experiment subsection 4.4.
Cross-Modal Alignment.
While the complete prototypes contribute diverse modality knowledge to address data heterogeneity and missing modalities at modality-specific and modality-shared representation levels, simply imputing missing data with zeros for the absent modality can introduce noise. We introduce an alignment loss between modality-specific representations from different modalities to alleviate noise arising from representations of missing modalities. We apply the alignment loss to the projected modality-specific representations. To achieve this goal, we utilize the $\ell_{2}$ distance and formulate the distance loss on two modality-specific representations as:
$\mathcal{L}_{CMA}=\sum_{i\in D_{i}}\|z^{\prime}_{Ai}-z^{\prime}_{Bi}\|_{2}^{2},$ (12)
where $z^{\prime}_{Ai}=g_{2}(z_{Ai})$, $z^{\prime}_{Bi}=g_{2}(z_{Bi})\in\mathbb{R}^{d}$ represent the projected modality-specific representations for modalities A and B, respectively. Our objective is to align the representation from the absent modality with that of the existing modality, thereby mitigating the negative impact in the missing modality. Finally, the overall training objective for each client is formulated as follows:
$\mathcal{L}=\mathcal{L}_{S}+\alpha_{reg}\mathcal{L}_{CMPR}+\alpha_{con}\mathcal{L}_{CMPC}+\alpha_{align}\mathcal{L}_{CMA}$ (13)
where $\alpha_{reg}$, $\alpha_{con}$ and $\alpha_{align}$ are hyper-parameters that can be tuned to control the importance of the $\mathcal{L}_{CMPR}$, $\mathcal{L}_{CMPC}$ and $\mathcal{L}_{CMA}$ terms, respectively. The overall training process for MFCPL is shown in Algorithm 1. The server broadcasts the complete prototypes to all clients in each communication round. In the local training phase, each client trains the learning model on private data using the overall training loss defined in Eq. (13).
Methods UCI-HAR (F1 score) Hateful Memes (AUC score) MELD (UAR score)
q = 0.5 q = 0.7 q = 0.8 q = 0.5 q = 0.7 q = 0.8 q = 0.5 q = 0.7 q = 0.8
FedAvg McMahan et al. (2017) 70.90 69.46 67.50 54.56 50.09 49.56 51.74 50.71 50.58
FedProx Li et al. (2020) 69.88 69.74 67.92 54.89 53.92 50.25 51.41 50.85 50.04
FedOpt Reddi et al. (2020) 74.79 66.17 56.45 54.62 51.61 49.58 50.96 49.28 48.27
MOON Li et al. (2021) 70.52 69.52 68.15 53.45 52.02 51.75 52.22 51.04 50.00
FedProto Tan et al. (2022b) 74.87 73.63 72.78 52.67 51.40 50.68 50.76 45.74 42.93
Ours (MFCPL) 77.44 75.61 75.19 56.30 55.69 55.04 53.73 52.24 51.92
Table 1: Performance comparison of our MFCPL with SOTA methods on UCI-HAR, Hateful Memes and MELD datasets with various modality missing rates. The best results are marked in bold.
4 Experiments
4.1 Experimental Setup
Datasets.
In this paper, we perform expeirments on three multimodal federated datasets to evaluate the effectiveness of MFCPL in various tasks.
  • UCI-HAR Anguita et al. (2013) dataset consists of smartphone sensory data from Accelerometer and Gyroscope sensors. This dataset involves observations from $30$ volunteers engaged in six distinct activities: walking, walking upstairs, walking downstairs, sitting, standing, and laying.
  • Hateful Memes Kiela et al. (2020) dataset comprises $10,000$ samples of image and text pairs organized into binary classes. The primary objective is to identify and detect hateful speech within memes.
  • MELD Poria et al. (2018) dataset is a multimodal conversational dataset for emotion recognition collected from Friends TV series, which contains about $13,000$ utterances from $1,433$ dialogues. Following the setting from Feng et al. (2023), we keep $4$ emotion labels with most samples which are neutral, sadness, happiness and anger.
Following Feng et al. (2023), we initialize $50$, $86$, and $105$ clients with the participation rates of $0.25$, $0.1$, and $0.1$ per communication round for Hateful Memes, MELD, and UCI-HAR datasets, respectively. To simulate the non-IID scenario, we apply the Dirichlet distribution Ferguson (1973) with $\beta=0.2$ to the Hateful Memes and UCI-HAR datasets. In contrast, the MELD dataset inherently exhibits a natural non-IID data distribution due to its organization by speakerIDs.
Model Architecture.
In the case of modalities such as image, text and audio, we follow Feng et al. (2023) to extract features for local training. This method involves deriving features instead of training from scratch using raw data, a choice made to alleviate the computational burden. Our model architecture is composed of three key modules: the encoder, the fusion layer, and the classifier. For audio, accelerometer, and gyroscope data, the encoder adopts a 1D Conv+RNN architecture. The 1D Conv includes three CNN layers followed by two MLP layers, while the RNN component incorporates a single GRU layer. In contrast, the encoder for images is built with two MLP layers, and for text, we employ an RNN architecture with a single GRU layer. The fusion layer is implemented as a self-attention fusion layer, and the classifier is subsequently followed by two MLP layers. The hidden size for GRU layer is set as $128$ for all datasets. Regarding the projection head, we utilize a single MLP layer with the default dimension set to $64$.
Implementation Details.
Following the methodology outlined in Feng et al. (2023), we set a uniform missing rate of $q=\{0.5,0.7,0.8\}$ for each modality based on Bernoulli distribution, creating a scenario that simulates the presence of missing modalities. The communication round is set to $200$, and the local training epoch is $1$ for all datasets. The default value for the temperature $\tau$ is $0.1$. The batch size is set to $16$, and we employ the SGD Robbins and Monro (1951) optimizer with a weight decay of $51e-$ 5, maintaining a consistent learning rate of $0.05$ across all datasets. We tunedhyperparameters $\alpha_{reg}$, $\alpha_{con}$, and $\alpha_{align}$ within the range of $\{0.1,3\}$ and configured with values yielding the best results for each dataset. For the Hateful Memes dataset, we use the AUC score as the evaluation metric. The UAR score serves as the metric for the MELD dataset, and the F1 score is employed as the metric for the UCI-HAR dataset. We repeat the experiment $5$ times and report mean values.
Baselines.
For evaluation, we compare our MFCPL with five baselines in MFL setting: (1) FedAvg McMahan et al. (2017), (2) FedProx Li et al. (2020), (3) FedOpt Reddi et al. (2020), (4) MOON Li et al. (2021), (5) FedProto Tan et al. (2022b) (with parameter averaging).
Figure 3: Illustrative example of average F1 score versus communication rounds on UCI-HAR dataset with $q=0.5$.
CMPR CMPC CMA UCI-HAR (F1 score) Hateful Memes (AUC score) MELD (UAR score)
q = 0.5 q = 0.7 q = 0.8 q = 0.5 q = 0.7 q = 0.8 q = 0.5 q = 0.7 q = 0.8
70.90 69.46 67.50 54.56 50.09 49.56 51.74 50.71 50.58
75.09 74.31 73.63 53.96 54.92 54.15 52.16 51.69 49.98
71.05 70.55 71.68 51.42 51.58 50.23 51.83 50.45 50.43
73.56 70.96 70.41 51.79 51.17 50.63 51.77 52.05 51.03
77.44 75.61 75.19 56.30 55.69 55.04 53.73 52.24 51.92
Table 2: Ablation study on key components of our MFCPL for three datasets under various modality missing rates.
4.2 Experimental Results
Performance Comparison.
Table 1 illustrates the performance comparison of our proposed MFCPL with other state-of-the-art methods on three multimodal federated datasets. Our method, MFCPL, consistently outperforms other baselines across various missing rates, highlighting its effectiveness in achieving better generalized global models and supporting clients with missing data. As the missing rate parameter $q$ increases, MFCPL maintains its performance, while other baselines exhibit a significant decrease in accuracy. Taking the UCI-HAR dataset with $q=0.5$ as an illustrative example, our method surpasses the best-performing baseline FedProto by a notable margin, showcasing an improvement of $2.57\%$. These results underscore the robustness and effectiveness of MFCPL in handling missing data scenarios across different client types and datasets. Figure 3 depics the performance curves (F1 score versus communication rounds) of our MFCPL and SOTA baselines, from which we can observe that MFCPL demonstrates a more stable convergence compared to other methods, particularly under non-IID and missing modality conditions.
Prototype UCI-HAR Hateful Memes MELD
Unimodal Prototypes 77.09 54.95 52.13
Complete Prototypes 77.44 56.30 53.73
Table 3: Performance analysis of MFCPL on utilizing complete prototypes and unimodal prototypes with $q=0.5$.
4.3 Comparison to Unimodal Prototypes
To assess the efficacy of our proposed complete prototypes, we conducted experiments comparing them with unimodal prototypes using $q=0.5$, as summarized in Table 3. The results demonstrate the superior performance of utilizing our proposed complete prototypes compared to the unimodal prototypes. Notably, our method exhibits a performance improvement of $1.6\%$ on the MELD dataset over the unimodal prototypes. This finding highlights the effectiveness of capturing diverse modality knowledge from complete prototypes, providing valuable guidance for the training process.
4.4 Ablation Studies
Effects of Key Components.
To comprehensively evaluate the impact of individual components on the performance of MFCPL, we conducted an ablation study by selectively removing each component: CMPR, CMPC, and CMA. The results are summarized in Table 2. Our findings indicate that leveraging all three components leads to the optimal performance across all three datasets, each characterized by distinct missing rates. Significantly, CMPC and CMA demonstrate a more pronounced impact compared to CMPR. This observation highlights the critical importance of ensuring alignment between complete prototypes and modality-specific representations, as well as the effectiveness of noise reduction in supporting clients with missing modalities. Additionally, we observed that the contribution of CMPR positively impacts the learning process, providing further insights into its role in enhancing overall performance.
Temperature parameter.
We further explore the overall performance under different values of $\tau$ with $q=0.5$, as depicted in Figure 3(a). Notably, we observe that a smaller temperature tends to enhance the training process more effectively than a larger one. It is worth noting that excessively high temperatures may exhibit a slightly negative effect on the overall performance.
Dimension of projection head.
We vary the dimension of the projection head $d$ and present the results in Figure 3(b). Our observations indicate that the MFCPL method exhibits robustness across different dimension sizes $d$. However, a noteworthy trend emerges: as we increase the dimension to larger sizes, there is a slight decrease in performance. This finding suggests that a smaller projection head dimension can minimize redundant information and enhance performance.
(a) Temperature $\tau$
(b) Dimension $d$
Figure 4: Performance analysis of MFCPL on three datasets under different values of temperature $\tau$ and projection head dimension $d$ with $q=0.5$.
5 Conclusion
In this paper, we propose MFCPL, a novel multimodal federated learning framework, breaking through the twin barriers of data heterogeneity and severely missing modalities. Our approach leverages a unique complete prototype methodology, incorporating three key components. We initiate the alignment of complete prototypes at two levels: modality-specific and modality-shared representations, thereby extracting rich semantic representation from multiple modalities. Additionally, we integrate cross-modal alignment to alleviate noise stemming from missing modalities, resulting in significant performance gains. Experiments show that our MFCPL consistently outperforms other state-of-the-art methods, achieving up to $7.69\%$, $5.48\%$ and $8.99\%$ improvements in performance for UCI-HAR, Hateful Memes, and MELD datasets, respectively when $80\%$ of modalities are missing. Thus, MFCPL paves the new way for robust multimodal learning in real-world settings with diverse and incomplete data.